Need help on solving these two problems, thanks =0)
1. -9x+20=-2x
2. 2x^2-5=3x

Answer by corrie21(4)

1. -9x+20=-2x
First, add 9x to -9x, these cancel out. Then add 9x to the -2x.
20=7x
to get the x by itself, divide the 7 into the 20.
20/7=x

2. 2x^2-5=3x
First, minus the 3x from both sides.
2x^2-3x-5=0
then factor out
(2x-5)(x+1)